# Progressive Binarization with Semi-Structured Pruning for LLMs

### v3

@@ Abstract
Large language models (LLMs) have achieved remarkable progress in natural language processing, but their high computational and memory costs hinder deployment on resource-constrained devices. Binarization, which reduces model weights to 1 bit, is a promising solution for efficient inference. However, binarized LLMs still exhibit redundancy that can be further compressed. Semi-structured pruning offers a favorable trade-off between model performance and hardware efficiency, but naively combining it with binarization often leads to severe performance degradation. To address this, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P), a novel post-training compression framework. We propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO) to jointly reduce pruning and binarization error. Additionally, we develop a Coarse-to-Fine Search (CFS) strategy to more effectively select pruning elements. Extensive experiments across multiple LLM families show that PBS2P consistently outperforms state-of-the-art binary post-training quantization methods in both perplexity and downstream accuracy. The code and models will be available at https://github.com/XIANGLONGYAN/PBS2P.
@@ 1 Introduction
Transformer-based large language models (LLMs) vaswani2017attention have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks. This success is largely driven by their massive scale, often comprising billions of parameters. For instance, the OPT (Open Pre-trained Transformer) series zhang2022opt includes models with up to 66 billion parameters, while the LLaMA family touvron2023llama1 features even larger models, such as LLaMA3-70B dubey2024llama3 . Although these models have substantially advanced the state-of-the-art (SOTA) in NLP, their increasingly computational and memory demands present significant challenges for deployment on the resource-constrained devices.
@@
The compression of LLMs can be broadly categorized into several approaches, including weight quantization lin2024awq ; frantar2022gptq , low-rank factorization zhang2023loraprune ; yuan2023asvd , network pruning sun2023simple ; frantar2023sparsegpt , and knowledge distillation zhong2024revisiting ; gu2023knowledge . Among these, binarization represents an extreme form of quantization, reducing model weights to 1 bit and drastically lowering memory consumption. Recent efforts primarily focus on post-training quantization (PTQ) huang2024billm ; li2024arb ; dong2024stbllm , which eliminates the need for backpropagation, thereby accelerating the binarization process and improving deployment efficiency. For example, BiLLM huang2024billm introduces a residual approximation strategy to enhance the performance of 1-bit LLMs, while ARB-LLM li2024arb adopts an alternating refinement approach to better align binarized weights with their full-precision counterparts. STBLLM dong2024stbllm further advances this line of work by compressing LLMs to sub-1-bit precision. Despite the aforementioned advancements, binarized LLMs continue to exhibit redundancy, indicating opportunities for additional compression.
@@
Pruning lecun1989optimal is a promising technique for further reducing redundancy in binarized models. However, conventional structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation often leads to severe performance degradation in LLMs. On the other hand, unstructured pruning dong2024pruner suffers from poor hardware compatibility and inefficient storage. Semi-structured pruning frantar2023sparsegpt ; sun2023simple ; dong2024stbllm has emerged as an effective compromise, significantly reducing redundancy while maintaining a balance between accuracy and hardware efficiency. Nonetheless, directly combining binarization with semi-structured pruning often leads to noticeable performance degradation, making it hard to maintain model accuracy. Another challenge is how to select pruning elements effectively to improve pruning efficiency while preserving accuracy.
@@
To address these challenges, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLMs, which achieves substantial model compression while maintaining strong performance (see Figure 1). We first propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively prunes a subset of elements at each step while jointly optimizing the binarized parameters. This strategy effectively reduces the cumulative error from pruning and binarization. To further enhance pruning efficiency, we develop a Coarse-to-Fine Search (CFS) algorithm for more accurate pruning element selection. In the coarse stage, pruning ratios are assigned to each layer based on its importance. In the fine stage, a Hessian-based metric is used to identify the most redundant elements, guided by the layer-specific pruning ratios. We further show that the Hessian-based metric minimizes the theoretical error increase under a second-order approximation. This makes it a provably optimal strategy for pruning.
@@
Extensive experiments demonstrate that PBS2P consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing binary PTQ methods on a wide range of evaluation benchmarks. As shown in Figure 1, on the WikiText-2 merity2016pointer dataset, PBS2P attains a perplexity of 6.20 on LLaMA-13B touvron2023llama1 with an average bit-width of just 0.8 bits, compared to 5.47 for the full-precision model. These results further highlight the effectiveness of PBS2P in significantly narrowing the performance gap between binarized and full-precision models.
@@ 2.1 LLM Quantization
Quantization compresses full-precision parameters into lower-bit representations, reducing both computation and storage demands. Current quantization methods for LLMs are mainly divided into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT liu2023llm ; chen2024db ; du2024bitdistiller integrates quantization during the training phase to enhance low-bit weight representations. However, due to the enormous parameter number, retraining becomes excessively expensive and inefficient for LLMs. PTQ, as it directly applies quantization to the model weights without retraining, making it faster and less resource-demanding. Recent methods, like ZeroQuant yao2022zeroquant and BRECQ li2021brecq , improve quantization accuracy by incorporating custom quantization blocks and group labels. While GPTQ frantar2022gptq and QuIP chee2024quip use second-order error compensation to reduce quantization errors.
@@ 2.2 Network Binarization
Binarization, as the most extreme form of quantization, reduces model parameters to a single bit (±1). Prominent methods, like Binary Weight Network (BWN) rastegari2016xnor , XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi , focus on binarizing the weights, with XNOR-Net rastegari2016xnor and Bi-Real Netliu2018bi also binarizing activations. In the context of LLM binarization, BitNet wang2023bitnet , OneBit (xu2024onebit, ), and BinaryMoS jo2024mixture adopt the QAT framework, while BiLLM huang2024billm , ARB-LLM li2024arb , and STBLLMdong2024stbllm use PTQ combined with residual approximation. Our work focuses on the binary PTQ method, achieving substantial improvements over existing SOTA binary PTQ methods.
@@ 2.3 Model Pruning
Pruning lecun1989optimal is a common technique for compressing neural networks by removing less important parameters, resulting in smaller and more efficient sparse models. In LLMs, pruning methods are typically categorized as structured, unstructured, or semi-structured. Structured pruning ma2023llm ; ashkboos2024slicegpt ; xia2023sheared ; an2024fluctuation removes entire components to enhance efficiency but often causes performance drops, requiring retraining. Unstructured pruning dong2024pruner eliminates individual weights based on importance, preserving performance even at high sparsity, though the irregular patterns are hardware-unfriendly. Semi-structured pruning offers a good trade-off, enforcing regular patterns like $N$:$M$ sparsity for hardware efficiency, as used in SparseGPT frantar2023sparsegpt , Wanda sun2023simple , and STBLLM dong2024stbllm . Our method adopts semi-structured $N$:$M$ sparsity to retain performance while ensuring hardware compatibility.
@@ 3.1 Preliminary
Binarization.    First, let’s briefly review the process of standard binarization rastegari2016xnor . Binarization begins by performing a row-wise redistribution on the full-precision matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ , resulting in a matrix $\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}$ with a mean of 0:
$\widetilde{\mathbf{W}}=\mathbf{W}-\mu,\quad\text{where}\,\mu=\frac{1}{m}\sum_{j=1}^{m}\mathbf{W}_{.j}\,.$ (1)
The binarization objective is to minimize the reconstruction error: $\arg\min_{\alpha,\mathbf{B}}\|\widetilde{\mathbf{W}}-\alpha\mathbf{B}\|_{F}^{2}$, where $\alpha\in\mathbb{R}^{n}$ is a row-wise scaling factor and $\mathbf{B}\in\{+1,-1\}^{n\times m}$ is the binary matrix. The optimal solutions for $\alpha$ and $\mathbf{B}$ are given by $\alpha=\frac{1}{m}\sum_{j=1}^{m}|\widetilde{\mathbf{W}}_{.j}|$ and $\mathbf{B}=\operatorname{sign}(\widetilde{\mathbf{W}})$, respectively huang2024billm .
N:M Sparsity.    $N$:$M$ sparsity is a semi-structured pruning scheme that balances model compression and hardware efficiency. Unlike unstructured pruning, which removes arbitrary weights and results in irregular sparsity patterns that are hard to accelerate, $N$:$M$ sparsity imposes a structured constraint by retaining exactly $N$ nonzero elements out of every $M$ consecutive weights frantar2023sparsegpt ; sun2023simple ; dong2024stbllm . A typical configuration, such as 2:4 sparsity, keeps only two weights out of every four, significantly reducing parameter count while preserving computational efficiency. This regular pattern allows for efficient execution on NVIDIA Ampere GPUs nvidia2020 , taking advantage of specialized hardware support. In our work, we adopt the $N$:$M$ sparsity scheme as a semi-structured pruning strategy to further eliminate redundancy within the binarized model. This combination enables more aggressive compression while maintaining hardware compatibility and inference efficiency.
@@ 3.2 Stepwise Semi-Structured Pruning with Binarization Optimization
We define the total error $\mathcal{L}$ for binarization and semi-structured pruning as follows:
$\mathcal{L}=||\mathbf{W}\mathbf{X}-(\mathbf{M}_{\text{p}}\odot\mathbf{\widehat{W}})\mathbf{X}||_{F}^{2},\quad\text{where}\,\widehat{\mathbf{W}}=\alpha\mathbf{B}+\mu\,.$ (2)
where $\mathbf{X}$ denotes the calibration data, $\mathbf{M}_{\text{p}}$ denotes the semi-structured pruning mask. Optimizing both the pruning mask $\mathbf{M}_{\text{p}}$ and the binarized parameters $\mathbf{\widehat{W}}$ simultaneously is an NP-hard problem blumensath2008iterative . Therefore, we adopt a greedy-style approximation. At each step, we prune a small portion of elements. Then, we update the binarized parameters to fit the current pruning ratio. This step-by-step strategy reduces the optimization difficulty. It distributes the optimization burden across multiple pruning stages. The selection of the semi-strucutred pruning mask $\mathbf{M}_{\text{p}}$ will be elaborated in Section 3.3.
Figure 3: Illustration of the $N$:$M$ sparsity pruning mask group, where the original pruning mask $\mathbf{M}_{\text{p}}$ is divided into $M$-$N$ sub masks.
N:M Sparsity Pruning Mask Group.
To enable progressive pruning, we decompose the original pruning mask $\mathbf{M}_{\text{p}}$ into a series of sub-masks, each representing a step of pruning. Under the $N$:$M$ sparsity constraint, exactly $N$ out of every $M$ elements are retained. Within each $M$-element group, the $M{-}N$ pruned elements are removed one by one across successive steps. Formally, the original pruning mask $\mathbf{M}_{\text{p}}$ can be expressed as the union of $M{-}N$ sub-masks:
$\mathbf{M}_{\text{p}}=\bigcup_{i=1}^{M-N}\mathbf{M}_{\text{p}}^{i}.$ (3)
Figure 3 illustrates this process. This decomposition supports progressive pruning, facilitating tighter integration with binarization in subsequent stages.
Stepwise Pruning with Binarization Optimization (SPBO).    After obtaining the group of pruning masks, we proceed with the stepwise pruning and binarization optimization process. We perform a total of $M{-}N$ pruning steps to achieve $N$:$M$ sparsity. After each step, we update the binarization parameters $\alpha$ and $\mu$ to better adapt to the current pruning state. At the $k$-th pruning step, we obtain the updated binary matrix $\mathbf{B}_{k}$ and full-precision matrix $\mathbf{W}_{k}$ by applying the $k$-th pruning sub-mask $\mathbf{M}_{\text{p}}^{k}$ to the previously pruned matrices $\mathbf{B}_{k-1}$ and $\mathbf{W}_{k-1}$:
@@
After the $k$-th pruning step, the binarization parameters $\alpha$ and $\mu$ are no longer optimal under the updated sparsity pattern. To minimize the error $\mathcal{L}$ in the current pruning state, these parameters must be re-estimated. Inspired by ARB-LLM li2024arb , a straightforward approach is to compute the partial derivatives of $\mathcal{L}$ with respect to $\alpha$ and $\mu$, and solve for the stationary points by setting them to zero. Specifically, we first update $\mu$ by setting $\partial\mathcal{L}/\partial\mu=0$ during the $k$-th pruning step:
@@
During the $k$-th pruning round, we can alternately update $\mu$ and $\alpha$ for $\tau$ steps to further reduce $\mathcal{L}$. We perform total $M$-$N$ steps of pruning until all pruning elements are removed. The detailed derivation of optimizing $\mu$ and $\alpha$ can be found in the supplementary material.

### v4

@@ Abstract
Large language models (LLMs) have achieved remarkable progress in natural language processing, but their high computational and memory costs hinder deployment on resource-constrained devices. Binarization represents the most extreme form of quantization, yet binarized models still contain redundancy that can be further removed. Pruning provides a natural way to eliminate such redundancy, but naïve combination with binarization often results in severe performance degradation. In this paper, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P), a novel post-training framework that seamlessly integrates binarization and semi-structured pruning. We first propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively introduces sparsity while optimizing binarization parameters to jointly reduce pruning and quantization error, yielding more stable and accurate compression. Additionally, we propose a Coarse-to-Fine Search (CFS) that first allocates pruning ratios and then refines element selection, further enhancing overall performance. Extensive experiments across multiple LLM families show that PBS2P consistently outperforms state-of-the-art (SOTA) binary post-training quantization methods in both perplexity and downstream accuracy. The code and models will be available at https://github.com/XIANGLONGYAN/PBS2P.
@@ 1 Introduction
Transformer-based large language models (LLMs) (Vaswani, 2017) have demonstrated remarkable performance across a wide range of natural language processing (NLP) tasks. This success is largely driven by their massive scale, often with billions of parameters. For instance, the OPT series (Zhang et al., 2022) includes models with up to 66 billion parameters, while the LLaMA family (Touvron et al., 2023a) features even larger models, such as LLaMA3-70B (Dubey et al., 2024). Although these models have substantially advanced the state of the art (SOTA) in NLP, their growing computational and memory demands present increasingly significant challenges for real-world deployment.
@@
The compression of LLMs can be broadly categorized into several approaches, including weight quantization (Lin et al., 2024; Frantar et al., 2023), low-rank factorization (Zhang et al., 2024; Yuan et al., 2023), network pruning (Sun et al., 2024; Frantar & Alistarh, 2023), and knowledge distillation (Zhong et al., 2024; Gu et al., 2024). Among these, binarization represents an extreme form of quantization, reducing model weights to 1 bit and drastically lowering memory consumption. Recent efforts primarily focus on post-training quantization (PTQ) (Huang et al., 2024; Li et al., 2025; Dong et al., 2025), which eliminates the need for backpropagation, thereby accelerating the binarization process and improving deployment efficiency. For example, BiLLM (Huang et al., 2024) introduces a residual approximation strategy to enhance the performance of 1-bit LLMs, while ARB-LLM (Li et al., 2025) adopts an alternating refinement approach to better align binarized weights with their full-precision counterparts. STBLLM (Dong et al., 2025) further advances this line of work by compressing LLMs to sub-1-bit precision. Despite the aforementioned advancements, binarized LLMs continue to exhibit redundancy, indicating opportunities for additional compression.
@@
Pruning (LeCun et al., 1989) is a promising technique for further reducing redundancy in binarized models. However, conventional structured pruning (Ma et al., 2023; Ashkboos et al., 2024; Xia et al., 2024; An et al., 2024) often leads to severe performance degradation in LLMs. On the other hand, unstructured pruning (Dong et al., 2024) suffers from poor hardware compatibility and inefficient storage. Semi-structured pruning (Frantar & Alistarh, 2023; Sun et al., 2024; Dong et al., 2025) has emerged as an effective compromise, significantly reducing redundancy while maintaining a balance between accuracy and hardware efficiency. Nonetheless, directly combining binarization with semi-structured pruning often leads to noticeable performance degradation, as the joint constraints on weight values and structure increase the difficulty of preserving model accuracy under extreme compression. This problem is further complicated by the need to select pruning elements carefully, since suboptimal choices may remove important weights and amplify quantization errors. Therefore, designing a unified framework that balances compression and accuracy retention remains challenging.
@@
To address these challenges, we propose Progressive Binarization with Semi-Structured Pruning (PBS2P) for LLMs, which achieves substantial model compression while maintaining strong performance (see Figure 1). We first propose Stepwise semi-structured Pruning with Binarization Optimization (SPBO), which progressively prunes a subset of elements at each step while jointly optimizing the binarized parameters. This strategy effectively reduces the cumulative error from pruning and binarization. To further enhance pruning efficiency, we develop a Coarse-to-Fine Search (CFS) algorithm for more accurate pruning element selection. In the coarse stage, pruning ratios are assigned to each layer based on its importance. In the fine stage, a Hessian-based metric is used to identify the most redundant elements, guided by the layer-specific pruning ratios. We further show that the Hessian-based metric minimizes the theoretical error increase under a second-order approximation. This makes it a provably optimal strategy for pruning. Together, SPBO and CFS form the core of PBS2P, enabling provably efficient compression with minimal accuracy degradation.
@@
Extensive experiments demonstrate that PBS2P consistently achieves SOTA performance across multiple LLM families, clearly surpassing existing binary PTQ methods on a wide range of evaluation benchmarks. As shown in Figure 1, on the WikiText-2 (Merity et al., 2017) dataset, PBS2P attains a perplexity of 6.20 on LLaMA-13B (Touvron et al., 2023a) with an average bit-width of just 0.8 bits, compared to 5.47 for the full-precision model. These results further highlight the effectiveness of PBS2P in significantly narrowing the performance gap between binarized and full-precision models.
@@ 2.1 LLM Quantization
Quantization compresses full-precision parameters into lower-bit representations, reducing both computation and storage demands. Current quantization methods for LLMs are mainly divided into Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). QAT (Liu et al., 2024; Chen et al., 2024; Du et al., 2024) integrates quantization during the training phase to enhance low-bit weight representations. However, due to the enormous parameter number, retraining becomes excessively expensive and inefficient for LLMs. PTQ, as it directly applies quantization to the model weights without retraining, making it faster and less resource-demanding. Recent methods, like ZeroQuant Yao et al. (2022) and BRECQ (Li et al., 2021), improve quantization accuracy by incorporating custom quantization blocks and group labels. While GPTQ (Frantar et al., 2023) and QuIP (Chee et al., 2024) use second-order error compensation to reduce quantization errors.
@@ 2.2 Network Binarization
Binarization, as the most extreme form of quantization, reduces model parameters to a single bit (±1). Prominent methods, like Binary Weight Network (BWN) (Rastegari et al., 2016), XNOR-Net (Rastegari et al., 2016) and Bi-Real Net(Liu et al., 2018), focus on binarizing the weights, with XNOR-Net (Rastegari et al., 2016) and Bi-Real Net(Liu et al., 2018) also binarizing activations. In the context of LLM binarization, BitNet (Wang et al., 2023), OneBit (Xu et al., 2024), and BinaryMoS (Jo et al., 2024) adopt the QAT framework, while BiLLM (Huang et al., 2024), ARB-LLM (Li et al., 2025), and STBLLM(Dong et al., 2025) use PTQ combined with residual approximation. Focusing on the PTQ setting, our method yields significant enhancements over prior SOTA binary PTQ techniques.
@@ 2.3 Model Pruning
Pruning (LeCun et al., 1989) is a common technique for compressing neural networks by removing less important parameters, resulting in smaller and more efficient sparse models. In LLMs, pruning methods are typically categorized as structured, unstructured, or semi-structured. Structured pruning (Ma et al., 2023; Ashkboos et al., 2024; Xia et al., 2024; An et al., 2024) removes entire components to enhance efficiency but often causes performance drops, requiring retraining. Unstructured pruning (Dong et al., 2024) eliminates individual weights based on importance, preserving performance even at high sparsity, though the irregular patterns are hardware-unfriendly. Semi-structured pruning offers a good trade-off, enforcing regular patterns like NN:$M$ sparsity for hardware efficiency, as used in SparseGPT (Frantar & Alistarh, 2023), Wanda (Sun et al., 2024), and STBLLM (Dong et al., 2025). We adopt NN:$M$ sparsity to balance performance and hardware efficiency.
@@ 3.1 Preliminary
Binarization.  We begin by revisiting the standard binarization process (Rastegari et al., 2016). The full-precision matrix $\mathbf{W}\in\mathbb{R}^{n\times m}$ is first row-wise normalized by subtracting the mean of each row, yielding a zero-centered matrix $\widetilde{\mathbf{W}}\in\mathbb{R}^{n\times m}$:
$\widetilde{\mathbf{W}}=\mathbf{W}-\mu,\quad\text{where}\quad\mu=\frac{1}{m}\sum_{j=1}^{m}\mathbf{W}_{.j}\,.$ (1)
[…]
N:M Sparsity.  NN:$M$ sparsity is a semi-structured pruning scheme that balances model compression and hardware efficiency. Unlike unstructured pruning, which removes arbitrary weights and results in irregular sparsity patterns that are hard to accelerate, NN:$M$ sparsity imposes a structured constraint by retaining exactly $N$ nonzero elements out of every $M$ consecutive weights (Frantar & Alistarh, 2023; Sun et al., 2024; Dong et al., 2025). A typical configuration, such as 2:4 sparsity, keeps only two weights out of every four, significantly reducing parameter count while preserving computational efficiency. This regular pattern allows for efficient execution on NVIDIA Ampere GPUs (Nvidia, 2020), taking advantage of specialized hardware support. In our work, we adopt the NN:$M$ sparsity scheme as a semi-structured pruning strategy to further eliminate redundancy within the binarized model. This enables stronger compression with hardware-friendly execution.
@@ 3.2 Stepwise Semi-Structured Pruning with Binarization Optimization
To facilitate the joint optimization of binarization and semi-structured pruning, we formulate the total compression error $\mathcal{L}$ as the objective function to be minimized:
[…]
where $\mathbf{X}$ denotes the calibration data, $\mathbf{M}_{\text{p}}$ denotes the semi-structured pruning mask. Optimizing both the pruning mask $\mathbf{M}_{\text{p}}$ and the binarized parameters $\mathbf{\widehat{W}}$ simultaneously is an NP-hard problem (Blumensath & Davies, 2008). Therefore, we adopt a greedy-style approximation. At each step, we prune a small portion of elements. Then, we update the binarized parameters to fit the current pruning ratio. This step-by-step strategy reduces the optimization difficulty. It distributes optimization across multiple pruning stages. Selection of the semi-structured mask $\mathbf{M}_{\text{p}}$ is detailed in Section 3.3.
Figure 3: Illustration of the NN:$M$ sparsity pruning mask group, where the original pruning mask $\mathbf{M}_{\text{p}}$ is divided into $M$-NN sub masks.
N: M Sparsity Pruning Mask Group. To support progressive pruning, we decompose the original pruning mask $\mathbf{M}_{\text{p}}$ into a sequence of sub-masks, each corresponding to one intermediate pruning step. Under the NN:$M$ sparsity constraint, exactly $N$ out of every $M$ elements are retained. Within each $M$-element group, the MNM{-}N pruned elements are removed sequentially across successive stages. Formally, the full pruning mask can be expressed as the union of MNM{-}N sub-masks: $\mathbf{M}_{\text{p}}=\bigcup_{i=1}^{M-N}\mathbf{M}_{\text{p}}^{i}$. This decomposition enables fine-grained progressive pruning and allows for a tighter and more seamless integration with binarization in subsequent stages, as illustrated in Figure 3.
Stepwise Pruning with Binarization Optimization (SPBO).  After obtaining the group of pruning masks, we proceed with the stepwise pruning and binarization optimization process. We perform a total of MNM{-}N pruning steps to achieve NN:$M$ sparsity. After each step, we update the binarization parameters $\alpha$ and $\mu$ to better adapt to the current pruning state. At the $k$-th pruning step, we obtain the updated binary matrix $\mathbf{B}_{k}$ and full-precision matrix $\mathbf{W}_{k}$ by applying the $k$-th pruning sub-mask $\mathbf{M}_{\text{p}}^{k}$ to the previously pruned matrices $\mathbf{B}_{k-1}$ and $\mathbf{W}_{k-1}$:
@@
After the $k$-th pruning step, the binarization parameters $\alpha$ and $\mu$ are no longer optimal under the updated sparsity pattern. To minimize the error $\mathcal{L}$ in the current pruning state, these parameters must be re-estimated. Inspired by Li et al. (2025), a straightforward approach is to compute the partial derivatives of $\mathcal{L}$ with respect to $\alpha$ and $\mu$, and solve for the stationary points by setting them to zero. Specifically, we first update $\mu$ by setting $\partial\mathcal{L}/\partial\mu=0$ during the $k$-th pruning step:
@@
During the $k$-th pruning round, we can alternately update $\mu$ and $\alpha$ for $\tau$ steps to further reduce $\mathcal{L}$. We perform total $M$-NN steps of pruning until all pruning elements are removed. The detailed derivation of optimizing $\mu$ and $\alpha$ can be found in the supplementary material.